OpenStudy (anonymous):

I'm confused so im in 8th grade and my math teacher assigned homework and said 2 point slope what would a 2 point slope look like im confused. Thanks:)

OpenStudy (astrophysics):

What's the question?

OpenStudy (astrophysics):

2 point slope, meaning a line

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

like on a graph i guess

OpenStudy (astrophysics):

 \[slope = \frac{ rise }{ run } = \frac{ \Delta y }{ \Delta x } = \frac{ y_2-y_1 }{ x_2-x_1 }\]